OpenStudy (anonymous):

How do I solve: p^2+2p+1=49 Please include steps. Thank you very much!

OpenStudy (anonymous):

\[p^2+2p+1=49\]

OpenStudy (campbell_st):

make it p^2 +2p - 48 = 0 factorise (p + 8)(p-6)=0 then p = -8, 6

razor99 (razor99):

yeah its coreect 100%

OpenStudy (anonymous):

always make the equation equal to zerp, which means you have to subtract 49 from bothsides and then factor and you'll get two values of x just like campbell_st did

razor99 (razor99):

@Delicia she is correct

razor99 (razor99):

i mean she/he

OpenStudy (anonymous):

Oh, I see. I forgot that I had to make the equation equal to 0, so I was stuck. Thanks for the help guys! :)